OpenStudy (anonymous):

Medal and Fan for answer. Which of these is a literal equation? ten minus six equals four three squared plus four equals ten plus three x squared minus seven equals y plus four three time eight minus ten equals three time negative two

OpenStudy (anonymous):

ten minus six equals four = 10 - 6 = 4 three squared plus four equals ten plus three = 3^2 + 4 = 10 + 3 x squared minus seven equals y plus four = x^2 - 7 = y + 4 three time eight minus ten equals three time negative two = 3 * 8 - 10 = 3 * -2

OpenStudy (anonymous):

Well the literary term for theequations would not be posted but if I had to choose it would be x squared minus seven equals y plus four

OpenStudy (anonymous):

Thank you both for the help, I appreciate it.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

your welcome.

OpenStudy (anonymous):

wrong!

OpenStudy (anonymous):

the first once: